OpenStudy (anonymous):

Find the derivative of the function. y = x / SQRT(x^6 + 4) y'=

OpenStudy (anonymous):

Do you know the quotient rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i don't know what it means

OpenStudy (anonymous):

So in other words the derivative of f(x) / g(x) , where f(x) is the numerator and g(x) is the denominator is ( f'(x)g(x) - f(x)g'(x) )/ (g(x))^2 So whats f(x) and g(x) in this case?

OpenStudy (anonymous):

f(x)=x/sqrt(x) g(x)=x^6 + 4 ?

OpenStudy (anonymous):

there are 3, right? so which one do you do?

OpenStudy (anonymous):

you're overthinking it, f(x) is the numerator, which is x and g(x) is sqrt (x^6 + 4)

OpenStudy (anonymous):

get that so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So we use the equation now right? so we have to find 3 things: f'(x), g'(x) and ((g(x))^2 You think you can do that?

OpenStudy (anonymous):

why g(x)^2?

OpenStudy (anonymous):

Isnt that part of the quotient rule?

OpenStudy (anonymous):

oh, ok.

OpenStudy (anonymous):

yeah so f'(x) = ? , g'(x) =? (g(x))^2 = ?

OpenStudy (anonymous):

f'(x)= 1 g'(x)= sqrt(6x^5) g(x)^2= x^6+4 ?

OpenStudy (anonymous):

pretty close, g'(x) = (sqrt (x^6 +4) )' right? so lets rewrite that as (x^6 + 4) ^ (1/2), im going to use the power rule? I'm not sure what you call it but im going to bring the 1/2 to the front, subtract one to get 1/2 (x^6 + 4) ^ (-1/2) * (x^6 +4)' , but we still need to use the chain rule on the inner part so we get (1/2(x^6+4)^(-1/2) ) (6x^5)

OpenStudy (anonymous):

thanks xNiker

OpenStudy (anonymous):

You got how to do it?

OpenStudy (anonymous):

i think so... it's (1/2(x^6+4)^(-1/2) ) (6x^5) right?

OpenStudy (anonymous):

Thats the derivative of g(x), after you got f' , g' and g^2, you gotta sub it all back in the quotient rule

OpenStudy (anonymous):

ok, that makes sense

OpenStudy (anonymous):

ok, wait. now i'm tripping up over the algebra \[\left( 1*1/2(x ^{6}+4)^{-1/2}*6x ^{5} \right)\div \left( x ^{6}+4 \right)\]

OpenStudy (anonymous):

\[y = x/ \sqrt(x^6 + 4)\] \[y' = ((1 * \sqrt(x^6+4)) - x * 1/2 (x^6 +4)^(-1/2) * (6x^5) ) / (x^6 +4)\]

OpenStudy (anonymous):

how do you simplify this?

OpenStudy (anonymous):

Does your teacher want you to simplfiy it? Idk